OpenStudy (anonymous):

how do you solve the solution for 2x-y=3, 3x+3y=5

OpenStudy (anonymous):

first use "2x-y=3" and u will find y=2x-3 then use it in "3x+3y=5" so that 3x+3 (2x-3)=5 then solve for x and y i think.

OpenStudy (rina.r):

so, u separate all the like terms first. 2x-3x-y-3y=5-3 then solve it.

OpenStudy (rina.r):

sorry collect all the like terms first.